### v1

# Data Distribution-based Curriculum Learning

Abstract
The order of training samples can have a significant impact on the performance of a classifier. Curriculum learning is a method of ordering training samples from easy to hard. This paper proposes the novel idea of a curriculum learning approach called Data Distribution-based Curriculum Learning (DDCL). DDCL uses the data distribution of a dataset to build a curriculum based on the order of samples. Two types of scoring methods known as DDCL (Density) and DDCL (Point) are used to score training samples thus determining their training order. DDCL (Density) uses the sample density to assign scores while DDCL (Point) utilises the Euclidean distance for scoring. We evaluate the proposed DDCL approach by conducting experiments on multiple datasets using a neural network, support vector machine and random forest classifier. Evaluation results show that the application of DDCL improves the average classification accuracy for all datasets compared to standard evaluation without any curriculum. Moreover, analysis of the error losses for a single training epoch reveals that convergence is faster when using DDCL over the no curriculum method.
Index Terms:
data distribution, curriculum learning, scoring method, classification, neural network, support vector machine, random forest.
I Introduction
Classification tasks in supervised machine learning use various techniques to create classifiers suited to the problem. Some of the widely used techniques are neural networks, support vector machines (SVM) and decision trees [1]. Neural networks are biologically inspired computing systems [2] that are widely used as classifiers due to their ability to learn and improve their performance through the use of data and experience. SVMs use a threshold to classify a data sample as belonging to one class or another class [3]. This is done by starting with the training data in a lower dimension and then moving to a higher dimension. The higher dimension data is then separated into groups that represent each class. Decision trees perform classification by using a tree-like structure made up of smaller decisions to make an overall decision for a given input. Multiple decision trees are often combined to create an ensemble classifier known as a Random Forest classifier [4].
The performance of these classifiers are dependent on the quality of data used and the robustness of training algorithms. Some of the factors that affect the quality of data are inaccurate, inconsistent, imbalanced, duplicate, missing and outlier samples in a dataset [5, 6]. Studies on the quality of data have shown that these properties may lead to a significant degradation in prediction performance and cause instability in learning due to high bias and/or high variance [5, 7].
To reduce the negative impact of these factors, the training algorithms applied to datasets are often designed to guide the learning model towards optimum performance [8, 9, 10]. Gradient descent is the most common method of optimising a neural network due to its fast convergence towards the minimum error [11]. Variants of gradient descent such as batch gradient descent, stochastic gradient descent (SGD) and mini batch gradient descent exist for use with specific problem scenarios. These problems may favour computation speed, data size or a balance of speed and size as frequently seen in practice [12]. Support vector machines are optimised by selecting a value for the C hyperparameter and the method of optimising its kernel varies according to the kernel used. The Radial Basis Function (RBF) kernel is commonly used in a SVM and it is optimised by selecting the gamma ($\gamma$) hyperparameter. Both of these hyperparameters can be found using several search algorithms including trees of Parzen estimators, particle swarm optimization and Bayesian optimisation. Recent findings have concluded that the first two algorithms provide better hyperparameter values with a lower execution time compared to Bayesian optimisation which has a very high computational cost [13]. In a random forest classifier, the number of decision trees (estimators) and the number of features required to split a tree are hyperparameters that are optimised [14].
Furthermore, the data distribution plays a major role in obtaining quality predictions from classifiers. When using the same data, a particular data distribution of training data can provide better results compared to another data distribution. Research has shown that the choice of distribution can cause an overall low bias or high bias for the same training data and also lead to poor performance without being apparent [15, 16]. This bias can have a significant impact on the model’s performance which can be critical in medical applications [15, 17].
Once the distribution of data is known, a decision can be made on how data samples are selected from the dataset for optimal results. This decision may include using a specific criteria for selecting samples in a particular order. The process of selectively choosing samples from a dataset as well as determining their order is known as creating a curriculum.
Using a curriculum for improving performance of a machine learning model is known as curriculum learning and it is typically applied on the order of training data samples. Curriculum learning enhances training of a model by starting with simple concepts and then gradually introducing difficult concepts as training progresses [18]. It is inspired by how humans are taught through curriculums in an education system by receiving basic education during childhood and then moving on to advanced education in adulthood [19, 20].
Curriculum learning has been applied to various problems achieving excellent results in the areas of image classification [21, 22], face recognition [23], visual attribute classification [24] and imbalanced data classification [25]. It has also been successfully used to improve performance of existing systems that use machine learning including unsupervised domain adaptation [26], transfer learning [27, 28] and reinforcement learning [29, 30].
In this paper, we propose a data distribution based curriculum learning approach for classification tasks. The proposed approach first determines the distribution of data and then information from the distribution is used to build a curriculum based on the order of samples. The approach is then evaluated on multiple datasets with classifiers based on three different types of widely used learning methods: neural networks, SVM and random forest classifier. The remainder of the paper is organised as follows. Section II provides a background on curriculum learning and related learning approaches as well as applications of curriculum learning to various problems. The proposed curriculum learning approach is discussed in Section III and Section IV outlines its experimental evaluation. Section V examines the experiment results while Section VI provides concluding remarks.
II Background
The concept of using a curriculum in machine learning was introduced by Bengio et al. in 2009 [18]. Their work defined curriculum learning as a training strategy where fewer simple concepts are presented in order at the beginning with greater complex concepts introduced at later stages. The authors conducted experiments on shape recognition and language modelling for scenarios with and without a curriculum. Experiment results showed that by using a curriculum, faster training and better convergence can be achieved.
Kumar et al. [31] built upon the concepts introduced by Bengio et al. by proposing self-paced learning (SPL), a training strategy that also suggested presenting training samples ordered from simple to complex. The main difference between SPL and curriculum learning was the process of defining the difficulty of a sample. In SPL, the current learning progress of a model was considered for selecting the next sample whereas curriculum learning relied on a pre-determined fixed curriculum. The authors conducted experiments that showed their proposed method outperforming the state of the art method.
Years later, Jiang et al. proposed a framework called self-paced curriculum learning (SPCL) that sought to unify the ideas of curriculum learning and self-paced learning [32]. They reasoned that both types of learning have drawbacks; The curriculum in curriculum learning is heavily reliant on the quality of fixed pre-determined prior knowledge and ignores feedback about the learner. In self-paced learning, the curriculum is dynamically determined to adjust to the learning pace of the learner but is prone to over-fitting due to being unable to handle prior knowledge. According to the authors, SPCL addresses these drawbacks by introducing a flexible way of including prior knowledge while also dynamically adjusting the curriculum based on feedback from the learner.
Hacohen et al. [28] expanded the definition of curriculum learning by describing it as consisting of two tasks with specific functions termed as the scoring function and the pacing function. The scoring function ranks the difficulty of the samples in a curriculum while the pacing function determines how often newer samples are presented to a model during training.
Curriculum learning has been applied to a variety of problems achieving promising results. Some of these problems include image classification [21, 22], face recognition [23], visual attribute classification [24] and imbalanced data classification [25]. Moreover, curriculum learning has also been successfully used to improve performance of systems where a curriculum is typically not considered. These include unsupervised domain adaptation [26], transfer learning [27, 28], bipedal walking for robots [30] and medical report generation [17].
Overall, we categorise the reviewed curriculum learning literature into three different groups: the applications of standard curriculum learning [24, 30, 17], variations of learning styles [27, 28, 23, 22] and research that makes use of the data density [21, 26, 25].
II-A Standard curriculum learning
Sarafianos et al. [24] applied curriculum learning to visual attribute classification by introducing a method that combined curriculum learning with multi-task learning. Their method performed end-to-end learning by providing a ConvNet with a complete image of a human without additional data to aid in classification. Then in multi-task learning, the tasks were split into groups using clustering. Curriculum learning was applied to the groups starting with the highest within-group cross-correlation and moving to the less correlated ones. Experiments performed on three datasets consisting of humans standing resulted in an increase in performance by up to 10%.
Tidd et al. [30] use curriculum learning to train deep reinforcement learning policies for bipedal walking of a robot over challenging terrain. The authors created a easy to hard curriculum using a three stage framework: In the first stage, guiding forces were applied to the joints and the base of the robot to start learning on easy terrain and gradually increased to difficult terrain. At the second stage when the terrain was most difficult, the guiding forces applied to the robot were slowly decreased. During the final stage, the magnitude of external random perturbations were increased to improve the robustness of the policy. Simulation experiments conducted by the authors demonstrated that a curriculum approach was effective in learning to walk for five types of terrain.
Ma et al. [17] applied curriculum learning to the medical report generation task. Their work introduced a framework capable of learning medical reports from limited medical data while reducing data bias. The framework learned in an easy to hard manner using a two step process where at first simple reports were used and then gradually complex reports consisting of rare and diverse medical abnormalities were attempted. This process effectively simulated the learning process of radiologists. The method was evaluated on two public datasets resulting in a boost in performance of the baselines.
Figure 1: Data Distribution-based Curriculum Learning
II-B Learning style variations
In [27], the Weinshall et al. used transfer learning to implement curriculum learning. They presented an approach where the curriculum was inferred through transfer learning from another network that was pre-trained on another task. A CNN with two architectures trained using curriculum learning was used to evaluate the proposed method on the CIFAR-100 and STL-10 datasets. Experiment results from their work concluded that with curriculum learning the convergence is faster during the beginning of training and improved generalisation is achieved when more difficult tasks are used.
Hacohen et al. applied curriculum learning to the training of deep networks [28]. They used transfer learning and bootstrapping as two different techniques for the easy to hard process which they termed the scoring function. With transfer learning, a ‘teacher’ network was trained on a large dataset and its prediction performance was used as the scoring function. In bootstrapping, a network was trained without a curriculum and its performance on the training data defined the scoring function. Then, the network was retrained from scratch using curriculum learning. For the pacing function, three approaches were used with all pacing functions having comparable performance based on experiments done by the authors. Experiments done on 6 test cases showed that using a curriculum provided high accuracy and better convergence with the amount of improvement ranging from small to large depending on the test case.
Huang et al. [23] propose a different approach to curriculum learning in the area of face recognition. Instead of using a traditional curriculum created by fixed ordering of samples with increasing difficulty, the authors introduce a loss function which incorporates adaptive curriculum learning. During the training process, samples are randomly selected for each mini-batch and the curriculum is created adaptively from the selected samples. Furthermore, the definition of hard samples is dynamic with a sample classified as hard during the start of training becoming easy towards the end. Overall, the loss function emphasizes easy samples at the start and hard samples later. The authors conducted experiments on benchmark data achieving improved performance over other methods.
Recent research has looked at improving the general curriculum learning process. Zhou et al. presented a curriculum learning method known as dynamics-optimized curriculum learning (DoCL) [22]. DoCL selected training samples at each step using weighted sampling based on the scores. The authors conducted experiments on more than 9 datasets achieving results that significantly improved the performance and efficiency compared to existing curriculum learning methods.
II-C Data density based curriculum learning
Our review of literature for curriculum learning has found few papers that utilise the density of the data. These works are presented in this section and our work on DDCL adds to this field of research.
Applying curriculum learning to large data sets has shown promising results. Researchers in [21] presented a method for training deep neural networks on large-scale weakly-supervised web images. Their training strategy utilised curriculum learning to effectively handle the large amount of noisy labels and data imbalance during the training process. They proposed a curriculum created by measuring the complexity of data using the density of its data. Experiments using the proposed training strategy resulted in state-of-the-art performance on benchmark data sets that included Webvision, ImageNet, Clothing-1M and Food-101.
Choi et al. [26] presented another method of using the density of data to generate the curriculum. Their method applied clustering on the data where higher density samples were considered simple and lower density samples were considered as complex. The proposed method was robust against false pseudo-labelled samples due to the use of a pseudo-labelling curriculum. They achieved state-of-the-art classification results on three benchmark datasets.
Curriculum learning has been applied to imbalanced data classification as well. Dynamic Curriculum Learning (DCL) proposed by Wang et.al [25] addressed the problem of requiring prior knowledge to train a system when using conventional techniques. DCL used a two-level curriculum scheduler made up of a sampling scheduler and a loss scheduler. The sampling scheduler found the best samples in a batch to train the model by dynamically managing the target data distributions from imbalanced to balanced and from easy to hard. The loss scheduler controlled the learning weights between classification loss and metric learning loss. Experiments achieved state-of-the-art performance on the CelebA face attribute dataset and the RAP pedestrian attribute dataset.
// Variables:
// Classes: $S$
// Data size: $N$
// $N_{s}$ data set for class $s$
// $|N_{s}|$ data size (cardinality)
// $h$ bandwidth
1 foreach $s\in S$ do
2       ${O_{s}}=\sum_{n=0}^{|N_{s}|}\min_{\mu_{s}\in S}(\parallel x_{n}-\mu_{s}\parallel^{2})$
3       $s=s\cup O_{s}$
4       $E_{s}=\varnothing$
5       foreach $x\in s$ do
6             $E_{s}=E_{s}\>\cup\parallel{O_{s}-x}\parallel$
7
8       end foreach
9      $D_{s}=\frac{1}{|N_{s}|h}\sum_{i=1}^{|N_{s}|}K(\frac{y-\hat{E}_{i}}{h})$
10       for $q\in Q^{s}$ do
11             if $|q|<3$ then
12                   $q=SMOTE(q)$
13             end if
14
15       end for
16      if Scoring type normal-$\leftarrow$ “Density” then
17            go to Algorithm 2
18       else if Scoring type normal-$\leftarrow$ “Point” then
19            go to Algorithm 3
20       end if
      Result: $T\leftarrow R_{s}$
21
22 end foreach
Algorithm 1 DDCL algorithm
III Proposed Curriculum Learning Approach
The curriculum learning approach proposed in this paper is based on the data distribution of a dataset. This is referred to as Data Distribution-based Curriculum Learning (DDCL) and involves multiple steps as visualised through Figure 1. It begins with dividing the data into groups based on their target classes. Then, the centroids for each group is calculated (Centroid Determination) and used to compute the Euclidean distance between each sample and its centroid. The data distribution for every class is determined next and utilised to divide the data for individual classes into quantiles (Q1, Q2, Q3). After quantile division, oversampling is optionally performed on quantiles where the number samples may be highly imbalanced compared to other quantiles. Lastly, the data is scored with either density or point scoring and rearranged for use in training.
The DDCL process is further detailed in Algorithm 1. In the first step of DDCL, the training data samples are grouped according to their classes ($S$) with the number of groups being equal to the number of unique classes in the data. Next, the grouped training data is used to calculate the centroids ($O_{s}$) for each class by performing clustering on each data group. Then, the centroids are used to compute the Euclidean distance from the relevant centroid ($E_{s}$) for each data sample ($x$). The third step determines the data distribution for a class ($D_{s}$) by plotting the normalised values of the Euclidean distances. Once the distribution of the training data is known, the data samples are divided into quantiles ($Q^{s}$). Each quantile ($q$) is then examined for the number of samples and the quantiles with the lowest number of samples are oversampled using SMOTE [33] if there are sufficient samples. SMOTE creates synthetic samples using the original data and different variants of it have been developed [34] since the initial method was proposed. The initial method of SMOTE is used in our proposed approach to address the potential lack of samples in a given quantile. Otherwise, no oversampling is applied due to insufficient samples. After the samples are divided into quantiles and optionally oversampled, each data sample is scored using either density or point scoring and rearranged accordingly ($R_{s}$). Finally, the rearranged training data ($T$) is ready to be passed to one of the learning methods for training and evaluation.
(a) Density scoring
(b) Point scoring
Figure 2: DDCL scoring methods
DDCL consists of two types of scoring methods: a sample density based method and a Euclidean distance based method. These are referred to as DDCL (Density) and DDCL (Point) respectively and are detailed in the following sections.
III-A Density based DDCL
Figure 1(a) illustrates the DDCL (Density) process where the scoring is applied to the quantiles based on the number of samples in the quantile. Quantiles with the greatest number of samples (Q3 - 86) are given a higher score while quantiles with the least number of samples (Q1 - 12) are given a lower score. Higher scores are considered simple while lower scores are considered difficult. This results in a final quantile order sorted from the highest to lowest density thus defining the curriculum to be used during training.
[…]
2 foreach $q\in Q^{s}$ do
3       $C_{s}=C_{s}\cup|q|$
4 end foreach
[…]
return $R_{s}$
Algorithm 2 DDCL Density
Algorithm 2 provides details of the DDCL (Density) scoring process. It shows that the process starts with determining the density of each quantile by taking their cardinality, $|q|$. Once the densities for all quantiles ($C_{s}$) are known, the training data is sorted according to their quantile’s density from highest to lowest. Finally, the rearranged data, $R_{s}$, is ready for use in training.
III-B Point based DDCL
On the other hand, Figure 1(b) illustrates how DDCL (Point) assigns a score to the individual data samples in the dataset by utilising their normalised Euclidean distance from centroid rather than examining quantiles. In this scoring method, data samples with the shorter Euclidean distances (P2, P5) are given the highest scores while samples with the longer Euclidean distances (P7, Pi) are assigned the lowest scores. As with DDCL (Density), higher scores in this point scoring method are considered simple while lower scores are considered difficult. Using DDCL (Point) results in a curriculum where the training samples are ordered based on their individual characteristic rather than as a group.
[…]
return $R_{s}$
Algorithm 3 DDCL Point
Algorithm 3 explains the DDCL (Point) scoring method and it shows that the data samples are sorted using the normalised Euclidean distance from centroid, $\hat{E_{s}}$. The sorting is done from the shortest distance to the longest regardless of which quantile a sample is assigned to. For example, consider a dataset with four quantiles with each quantile having at least five data samples. The first sample with the lowest $\hat{E_{s}}$ value may belong to quantile 1. Then, the sample with the second lowest $\hat{E_{s}}$ could come from quantile 4. The third lowest $\hat{E_{s}}$ value can be a sample from quantile 2. Next, the fourth lowest $\hat{E_{s}}$ may come from a sample in quantile 3. This method of sorting then repeats until $\hat{E_{s}}$ values from all data samples are accounted for with the resulting order of quantiles and samples ($R_{s}$) becoming randomised in contrast to DDCL (Density).
IV Experiments
DDCL is evaluated by performing classification tasks on 7 datasets obtained from the UCI Machine Learning Repository [35]. Table I lists the datasets used for the experiments. Three types of classifiers are used for evaluating DDCL; neural network, SVM and random forest.
[b] Dataset Data Instances Attributes Class Type Breast Cancer Diagnostic 569 31 Binary Cancer 457 9 Binary Haberman 306 3 Binary Liver Disorder 346 6 Binary Pima Indians Diabetes 768 8 Binary New-Thyroid 215 5 Multi-class Diabetes 1301 86556 46 Multi-class
TABLE I: Datasets used for evaluating DDCL.
  • 1
    Preprocessed version found in [36].
The neural network architecture used for evaluating DDCL consists of one input layer, a variable number of hidden layers and one output layer. The size of the input layer is equal to the number of attributes in the dataset and varies depending on the selected dataset. Similarly, the number of hidden layers changes based on the dataset and is a hyperparameter that can be tuned. Bayesian optimisation is used on the model to tune the hidden layer values. Finally, the size of the output layer can be configured for binary or multi-class classification. The SVM classifier for evaluating DDCL uses a C value of 1.0 with a RBF kernel. The gamma ($\gamma$) value of the RBF kernel scales according to the dataset used. A RBF kernel is chosen due to its good generalisation ability and robustness to input noise. The random forest classifier in our evaluation of DDCL uses 100 estimators and a value of 2 features for splitting a node.
During the neural network experiments, the selected dataset is split into training, validation and testing subsets. 20% of the data is used for validation, 10% for testing and the remainder for training. The neural network model is then trained on the training subset for 200 epochs and tested on the testing subset. Likewise in the SVM and random forest experiments, the selected dataset is split into training and testing subsets with 70% and 30% of the data being used respectively. The SVM and random forest classifiers are trained using the training subset of the data and tested on the testing subset.
For each dataset and classifier, 30 experiments runs are performed using three test scenarios: No Curriculum, DDCL (Density) and DDCL (Point). The first No Curriculum test scenario takes the input data and uses it to train a neural network, SVM and random forest classifier without altering the order of the training data. After training, classification experiments are conducted on the testing subset of data. In the second DDCL (Density) test scenario, the input data is processed using the steps outlined in the DDCL method and scored by density. Classification is then performed on the testing subset as stated earlier. The DDCL (Point) test scenario processes the input data through DDCL with the point scoring approach being used for the scoring step. Experiments for classification are likewise performed on the testing subset.
TABLE II: Neural Network classification results.
TABLE III: SVM classification results.
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 91.304 100.000 96.232 $\pm$2.442
DDCL (Density) 89.130 100.000 96.594 $\pm$2.673
DDCL (Point) 86.957 100.000 97.319 $\pm$2.953
Cancer No Curriculum 86.486 100.000 94.955 $\pm$3.993
DDCL (Density) 91.667 100.000 96.852 $\pm$2.658
DDCL (Point) 86.111 100.000 96.574 $\pm$3.100
Haberman No Curriculum 45.833 79.167 65.833 $\pm$8.898
DDCL (Density) 50.000 87.500 69.167 $\pm$10.353
DDCL (Point) 41.667 87.500 67.222 $\pm$9.424
Liver Disorder No Curriculum 57.143 85.714 69.048 $\pm$6.477
DDCL (Density) 46.429 85.714 68.214 $\pm$9.052
DDCL (Point) 57.143 82.143 69.524 $\pm$6.098
Pima Indians Diabetes No Curriculum 54.098 86.885 71.803 $\pm$6.658
DDCL (Density) 55.738 85.246 72.732 $\pm$5.933
DDCL (Point) 57.377 80.328 71.803 $\pm$5.460
New-Thyroid No Curriculum 82.353 100.000 93.137 $\pm$5.280
DDCL (Density) 82.353 100.000 95.490 $\pm$5.409
DDCL (Point) 82.353 100.000 94.118 $\pm$5.683
Diabetes 130 No Curriculum 53.235 54.694 54.029 $\pm$0.518
DDCL (Density) 56.148 56.798 56.579 $\pm$0.246
DDCL (Point) 54.400 56.278 55.139 $\pm$0.696
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 94.737 99.415 97.700 $\pm$1.188
DDCL (Density) 95.322 99.415 97.758 $\pm$1.025
DDCL (Point) 94.737 99.415 97.719 $\pm$1.239
Cancer No Curriculum 94.928 99.275 96.715 $\pm$1.134
DDCL (Density) 94.118 99.265 96.985 $\pm$1.145
DDCL (Point) 92.647 99.265 96.225 $\pm$1.562
Haberman No Curriculum 65.217 78.261 72.609 $\pm$4.051
DDCL (Density) 65.217 83.696 73.841 $\pm$4.836
DDCL (Point) 67.391 79.348 73.080 $\pm$3.329
Liver Disorder No Curriculum 44.231 76.923 63.750 $\pm$7.264
DDCL (Density) 49.038 75.000 66.699 $\pm$5.515
DDCL (Point) 55.769 75.962 68.558 $\pm$4.783
Pima Indians Diabetes No Curriculum 71.429 80.519 76.046 $\pm$2.500
DDCL (Density) 70.996 82.251 77.128 $\pm$2.991
DDCL (Point) 72.294 82.684 76.508 $\pm$2.514
New-Thyroid No Curriculum 92.308 100.000 97.128 $\pm$2.472
DDCL (Density) 92.308 100.000 97.282 $\pm$1.892
DDCL (Point) 92.308 100.000 96.872 $\pm$1.563
Diabetes 130 No Curriculum 64.285 65.487 64.948 $\pm$0.302
DDCL (Density) 64.220 65.457 64.977 $\pm$0.298
DDCL (Point) 64.301 65.469 65.052 $\pm$0.266
TABLE III: SVM classification results.
TABLE IV: Random Forest classification results.
Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
Breast Cancer Diagnostic No Curriculum 94.152 98.246 95.945 $\pm$1.034
DDCL (Density) 92.398 98.246 95.653 $\pm$1.581
DDCL (Point) 91.228 98.246 95.945 $\pm$1.517
Cancer No Curriculum 92.754 100.000 97.150 $\pm$1.374
DDCL (Density) 94.853 100.000 96.691 $\pm$1.298
DDCL (Point) 94.118 100.000 97.206 $\pm$1.361
Haberman No Curriculum 58.696 76.087 68.261 $\pm$4.100
DDCL (Density) 65.217 79.348 70.145 $\pm$3.825
DDCL (Point) 57.609 77.174 69.746 $\pm$4.367
Liver Disorder No Curriculum 59.615 78.846 70.417 $\pm$5.104
DDCL (Density) 67.308 77.885 72.917 $\pm$2.709
DDCL (Point) 60.577 78.846 70.064 $\pm$4.077
Pima Indians Diabetes No Curriculum 71.861 79.654 75.483 $\pm$1.985
DDCL (Density) 70.996 80.087 75.339 $\pm$2.200
DDCL (Point) 65.368 81.818 76.003 $\pm$3.461
New-Thyroid No Curriculum 95.385 100.000 97.949 $\pm$1.214
DDCL (Density) 92.308 100.000 98.000 $\pm$1.911
DDCL (Point) 95.385 100.000 97.897 $\pm$1.512
Diabetes 130 No Curriculum 65.830 66.658 66.218 $\pm$0.228
DDCL (Density) 65.626 66.948 66.242 $\pm$0.309
DDCL (Point) 65.542 66.805 66.199 $\pm$0.293
The results of all three test scenarios using neural networks, SVM and random forest classifiers are given in Table III, Table III and Table IV respectively.
Figure 3: Precision-Recall curves for binary classification datasets and confusion matrix for multi-class datasets.
V Discussion
Observations on the learning methods show that the classification performance across the datasets is always increased whenever DDCL is applied. This is true regardless of the learning method used as shown in Table III, Table III and Table IV.
Figure 4: Error loss per epoch for neural network classifier.
Using the neural network classifier provides the highest best results for the Haberman, Liver Disorder and Pima Indians Diabetes datasets while the SVM method achieves the highest average results on the Breast Cancer Diagnostic, Haberman and Pima Indians Diabetes data. The random forest based classifier yields the highest average result for the Cancer, Liver Disorder, New-Thyroid and Diabetes 130 datasets. Notably, using DDCL with SVM and random forest approaches results in significant increases in average accuracy for Haberman and Liver Disorder data respectively compared to other learning methods. Figure 3 shows the precision-recall curves for the best result of each binary classification dataset when using the neural network method. For multi-class data, the confusion matrix is provided in place of the precision-recall curves.
The highest average result for Breast Cancer Diagnostic was achieved using SVM with density-based DDCL while the largest improvement (1.087%) was using the neural network with point-based DDCL. For the Cancer dataset, using random forest with point-based DDCL provided the highest average accuracy and using the neural network with density-based DDCL gave the most improved accuracy (1.897%). Using SVM with density-based DDCL resulted in the highest average accuracy when evaluated on the Haberman data. The largest improvement for this dataset (3.334%) was observed when using density-based DDCL with the neural network. The highest average result for Liver Disorder was achieved using random forest with density-based DDCL while the SVM gave the largest average accuracy improvement at 4.808%. For the Pima Indians Diabetes dataset, the highest average accuracy was obtained using SVM with density-based DDCL which was also the most improved accuracy (1.082%).
In multi-class classification, using random forest with density-based DDCL resulted in the highest average accuracy on the New-Thyroid data and the largest improvement for this dataset came from density-based DDCL applied to the neural network. For the Diabetes 130 data, density-based DDCL with random forest provided the highest average accuracy while the largest improvement in accuracy (2.550%) was obtained through the neural network classifier utilising density-based DDCL.
The experiment results demonstrate that using DDCL to order the training data for a given learning method leads to improvements in classification accuracy. This can be seen in the performance of all datasets tested with either the density or point-based DDCL approach having increased performance compared to the no curriculum approach. Figure 4 shows the error loss per epoch plots when using the neural network classifier for each dataset except for Diabetes 130 which is excluded since its average accuracy is not significant [36]. The plots show that using DDCL leads to better reduction in the error loss.
In addition, using DDCL to order the training data results in faster convergence. We arrived at this conclusion by examining the error loss for the first five epochs of each dataset using batch gradient descent. Figure 5 shows the error loss plots for each dataset. It can be seen that the overall losses for DDCL (Density) and DDCL (Point) are reduced faster compared to the no curriculum approach thus leading to faster convergence towards the minimum. This is consistent with the findings of the error loss against epoch plots in Figure 4 where the error loss towards the end is lower for DDCL (Density) and DDCL (Point).
Figure 5: Error loss per epoch with batch gradient descent for neural network classifier.
VI Conclusion
This paper proposed a curriculum learning approach known as Data Distribution-based Curriculum Learning (DDCL). It used the data distribution of a dataset to build a curriculum based on the order of samples. Two types of scoring methods known as DDCL (Density) and DDCL (Point) were used to score samples thus determining their training order. DDCL (Density) used the sample density to assign scores while DDCL (Point) utilised the Euclidean distance for scoring.
Experiments were conducted on multiple datasets using a neural network, SVM and random forest classifier to evaluate DDCL. Results showed that while performance varied across the datasets depending on the classifier used, the application of DDCL resulted in significant increases in the average accuracy for most datasets compared to not using any curriculum. Furthermore, analysis of the error losses for five training epochs using batch gradient descent reveals that convergence is faster when using DDCL with either of the scoring methods over the no curriculum method.
The current DDCL approach uses only two types of scoring methods and is limited to using a single scoring method at a time. Moreover, the proposed approach uses a fixed curriculum that is pre-determined before the start of training and does not take into account the current training progress.
Future work will explore the creation of additional scoring methods to determine their impact on the training performance and investigate the application of ensemble learning using each scoring method. In addition, self-paced learning concepts will be incorporated into DDCL in order to dynamically determine the curriculum based on feedback from the learner. The proposed approach can also be adapted to image and text processing applications.

### v2

@@
-Data Distribution-based Curriculum Learning
+
 Abstract
-The order of training samples can have a significant impact on the performance of a classifier.
-Curriculum learning is a method of ordering training samples from easy to hard.
-This paper proposes the novel idea of a curriculum learning approach called Data Distribution-based Curriculum Learning (DDCL).
-DDCL uses the data distribution of a dataset to build a curriculum based on the order of samples.
-Two types of scoring methods known as DDCL (Density) and DDCL (Point) are used to score training samples thus determining their training order.
-DDCL (Density) uses the sample density to assign scores while DDCL (Point) utilises the Euclidean distance for scoring.
-We evaluate the proposed DDCL approach by conducting experiments on multiple datasets using a neural network, support vector machine and random forest classifier.
-Evaluation results show that the application of DDCL improves the average classification accuracy for all datasets compared to standard evaluation without any curriculum.
-Moreover, analysis of the error losses for a single training epoch reveals that convergence is faster when using DDCL over the no curriculum method.
+The order of training samples can have a significant impact on a model’s performance.
+Curriculum learning is an approach for gradually training a model by ordering samples from ‘easy’ to ‘hard’.
+This paper proposes the novel idea of a curriculum learning strategy called Data Distribution-based Curriculum Learning (DDCL).
+DDCL uses the inherent data distribution of a dataset to build a curriculum based on the order of samples.
+Our proposed approach is innovative as it incorporates two distinct scoring methods known as DDCL-Density and DDCL-Point to determine the order of training samples.
+The DDCL-Density method assigns scores based on the density of samples favoring denser regions that can make initial learning easier.
+Conversely, DDCL-Point utilizes the Euclidean distance from the centroid of the dataset as a reference point to score samples providing an alternative perspective on sample difficulty.
+We evaluate the proposed DDCL approach by conducting experiments across various classifiers using a diverse set of small to medium-sized medical datasets.
+Results show that DDCL improves the classification accuracy, achieving increases ranging from 2% to 10% compared to baseline methods and other state-of-the-art techniques.
+Moreover, analysis of the error losses for a single training epoch reveals that DDCL not only improves accuracy but also increases the convergence rate, underlining its potential for more efficient training.
+The findings suggest that DDCL can specifically be of benefit to medical applications where data is often limited and indicate promising directions for future research in domains that involve limited datasets.
 Index Terms:
-data distribution, curriculum learning, scoring method, classification, neural network, support vector machine, random forest.
+classification, curriculum learning, data distribution, machine learning, neural network, random forest, support vector machine
+[…]
+=-21pt
 I Introduction
@@
 To reduce the negative impact of these factors, the training algorithms applied to datasets are often designed to guide the learning model towards optimum performance [8, 9, 10].
-Gradient descent is the most common method of optimising a neural network due to its fast convergence towards the minimum error [11].
-Variants of gradient descent such as batch gradient descent, stochastic gradient descent (SGD) and mini batch gradient descent exist for use with specific problem scenarios.
-These problems may favour computation speed, data size or a balance of speed and size as frequently seen in practice [12].
-Support vector machines are optimised by selecting a value for the C hyperparameter and the method of optimising its kernel varies according to the kernel used.
-The Radial Basis Function (RBF) kernel is commonly used in a SVM and it is optimised by selecting the gamma ($\gamma$) hyperparameter.
-Both of these hyperparameters can be found using several search algorithms including trees of Parzen estimators, particle swarm optimization and Bayesian optimisation.
-Recent findings have concluded that the first two algorithms provide better hyperparameter values with a lower execution time compared to Bayesian optimisation which has a very high computational cost [13].
-In a random forest classifier, the number of decision trees (estimators) and the number of features required to split a tree are hyperparameters that are optimised [14].
+Guidance for smaller datasets is particularly important since their limited size cannot provide the sample diversity present in larger datasets [11].
+Gradient descent is the most common method of optimizing a neural network due to its fast convergence towards the minimum error [12].
+Variants of gradient descent such as batch gradient descent, stochastic gradient descent (SGD) and mini-batch gradient descent exist for use with specific problem scenarios [13].
+These problems may favor computation speed, data size or a balance of speed and size as frequently seen in practice [14].
+Support vector machines are optimized by selecting a value for the C hyperparameter and the method of optimizing its kernel varies according to the kernel used.
+The Radial Basis Function (RBF) kernel is commonly used in a SVM and it is optimized by selecting the gamma ($\gamma$) hyperparameter.
+Both of these hyperparameters can be found using several search algorithms including trees of Parzen estimators, particle swarm optimization and Bayesian optimization.
+Recent findings have concluded that the first two algorithms provide better hyperparameter values with a lower execution time compared to Bayesian optimization which has a very high computational cost [15].
+In a random forest classifier, the number of decision trees (estimators) and the number of features required to split a tree are hyperparameters that are optimized [16].
+The number of estimators for a specific problem is carefully selected since it corresponds with the computation time required.
 Furthermore, the data distribution plays a major role in obtaining quality predictions from classifiers.
 When using the same data, a particular data distribution of training data can provide better results compared to another data distribution.
-Research has shown that the choice of distribution can cause an overall low bias or high bias for the same training data and also lead to poor performance without being apparent [15, 16].
-This bias can have a significant impact on the model’s performance which can be critical in medical applications [15, 17].
+Research has shown that the choice of distribution can cause bias in the training data due to under-representation of the minority class [17].
+This bias can have a significant impact on a model’s accuracy and precision which can be critical for medical applications where training data can be limited [18, 19, 20].
 Once the distribution of data is known, a decision can be made on how data samples are selected from the dataset for optimal results.
@@
 The process of selectively choosing samples from a dataset as well as determining their order is known as creating a curriculum.
-Using a curriculum for improving performance of a machine learning model is known as curriculum learning and it is typically applied on the order of training data samples.
-Curriculum learning enhances training of a model by starting with simple concepts and then gradually introducing difficult concepts as training progresses [18].
+Using a curriculum for improving performance of a machine learning model is known as curriculum learning [21] and it is typically applied on the order of training data samples.
+Curriculum learning enhances training of a model by starting with simple concepts and then gradually introducing difficult concepts as training progresses [21, 22].
 It is inspired by how humans are taught through curriculums in an education system by receiving basic education during childhood and then moving on to advanced education in adulthood [23, 24].
 Curriculum learning has been applied to various problems achieving excellent results in the areas of image classification [25, 26], face recognition [27], visual attribute classification [28] and imbalanced data classification [29].
-It has also been successfully used to improve performance of existing systems that use machine learning including unsupervised domain adaptation [26], transfer learning [27, 28] and reinforcement learning [29, 30].
-In this paper, we propose a data distribution based curriculum learning approach for classification tasks.
+The effectiveness of the ‘easy’ to ‘hard’ strategy, where simpler examples are presented before more complex ones, has been well-documented in literature [30, 31].
+However, rival approaches offer different strategies by focusing on sample difficulty and order during training.
+Self-paced learning (SPL) [32], a training strategy that also suggested presenting training samples ordered from simple to complex, built upon the concepts introduced in curriculum learning by altering the process of defining the difficulty of a sample.
+In SPL, the current learning progress of a model was considered for selecting the next sample whereas curriculum learning relied on a pre-determined fixed curriculum.
+The concepts of SPL were used years later in a framework called self-paced curriculum learning (SPCL) [33].
+SPCL sought to unify the ideas of curriculum learning and SPL.
+The authors of the study reasoned that both types of learning have drawbacks; The curriculum in curriculum learning is heavily reliant on the quality of fixed pre-determined prior knowledge and ignores feedback about the learner.
+In SPL, the curriculum is dynamically determined to adjust to the learning pace of the learner but is prone to over-fitting due to being unable to handle prior knowledge.
+According to the authors, SPCL addressed these drawbacks by introducing a flexible way of including prior knowledge while also dynamically adjusting the curriculum based on feedback from the learner.
+Despite this, curriculum learning has been successfully used to improve performance of existing systems that use machine learning including unsupervised domain adaptation [34], transfer learning [35, 36] and reinforcement learning [30, 37].
+In this paper, we propose a data distribution based curriculum learning approach for classification tasks on small to medium-sized medical datasets.
 The proposed approach first determines the distribution of data and then information from the distribution is used to build a curriculum based on the order of samples.
 The approach is then evaluated on multiple datasets with classifiers based on three different types of widely used learning methods: neural networks, SVM and random forest classifier.
-The remainder of the paper is organised as follows.
-Section II provides a background on curriculum learning and related learning approaches as well as applications of curriculum learning to various problems.
+The remainder of the paper is organized as follows.
+Section II provides a background on curriculum learning as well as applications of curriculum learning to various problems.
 The proposed curriculum learning approach is discussed in Section III and Section IV outlines its experimental evaluation.
-Section V examines the experiment results while Section VI provides concluding remarks.
+Section V examines the experiment results whereas Section VI provides concluding remarks.
 II Background
@@
 Experiment results showed that by using a curriculum, faster training and better convergence can be achieved.
-Kumar et al.
-[31] built upon the concepts introduced by Bengio et al.
-by proposing self-paced learning (SPL), a training strategy that also suggested presenting training samples ordered from simple to complex.
-The main difference between SPL and curriculum learning was the process of defining the difficulty of a sample.
-In SPL, the current learning progress of a model was considered for selecting the next sample whereas curriculum learning relied on a pre-determined fixed curriculum.
-The authors conducted experiments that showed their proposed method outperforming the state of the art method.
-Years later, Jiang et al.
-proposed a framework called self-paced curriculum learning (SPCL) that sought to unify the ideas of curriculum learning and self-paced learning [32].
-They reasoned that both types of learning have drawbacks; The curriculum in curriculum learning is heavily reliant on the quality of fixed pre-determined prior knowledge and ignores feedback about the learner.
-In self-paced learning, the curriculum is dynamically determined to adjust to the learning pace of the learner but is prone to over-fitting due to being unable to handle prior knowledge.
-According to the authors, SPCL addresses these drawbacks by introducing a flexible way of including prior knowledge while also dynamically adjusting the curriculum based on feedback from the learner.
 Hacohen et al.
-[28] expanded the definition of curriculum learning by describing it as consisting of two tasks with specific functions termed as the scoring function and the pacing function.
-The scoring function ranks the difficulty of the samples in a curriculum while the pacing function determines how often newer samples are presented to a model during training.
+[36] expanded the definition of curriculum learning by describing it as consisting of two tasks with specific functions, termed the scoring function and the pacing function.
+The scoring function ranks the difficulty of the samples in a curriculum whereas the pacing function determines how often newer samples are presented to a model during training.
 Curriculum learning has been applied to a variety of problems achieving promising results.
@@
 These include unsupervised domain adaptation [34], transfer learning [35, 36], bipedal walking for robots [37] and medical report generation [19].
-Overall, we categorise the reviewed curriculum learning literature into three different groups: the applications of standard curriculum learning [24, 30, 17], variations of learning styles [27, 28, 23, 22] and research that makes use of the data density [21, 26, 25].
+Overall, we categorize the reviewed curriculum learning literature into three different groups: the applications of standard curriculum learning [28, 37, 19], variations of learning styles [35, 36, 27, 26] and research that makes use of the data density [25, 34, 29].
 II-A Standard curriculum learning
@@
 [28] applied curriculum learning to visual attribute classification by introducing a method that combined curriculum learning with multi-task learning.
-Their method performed end-to-end learning by providing a ConvNet with a complete image of a human without additional data to aid in classification.
+Their method performed end-to-end learning by providing a convolutional neural network (CNN) with a complete image of a human without additional data to aid in classification.
 Then in multi-task learning, the tasks were split into groups using clustering.
@@
 [37] use curriculum learning to train deep reinforcement learning policies for bipedal walking of a robot over challenging terrain.
-The authors created a easy to hard curriculum using a three stage framework: In the first stage, guiding forces were applied to the joints and the base of the robot to start learning on easy terrain and gradually increased to difficult terrain.
+The authors created an easy to hard curriculum using a three stage framework: In the first stage, guiding forces were applied to the joints and the base of the robot to start learning on easy terrain and gradually increased to difficult terrain.
 At the second stage when the terrain was most difficult, the guiding forces applied to the robot were slowly decreased.
@@
 Simulation experiments conducted by the authors demonstrated that a curriculum approach was effective in learning to walk for five types of terrain.
+Figure 1: Data Distribution-based Curriculum Learning
 Ma et al.
@@
 The method was evaluated on two public datasets resulting in a boost in performance of the baselines.
-Figure 1: Data Distribution-based Curriculum Learning
 II-B Learning style variations
@@
 They presented an approach where the curriculum was inferred through transfer learning from another network that was pre-trained on another task.
-A CNN with two architectures trained using curriculum learning was used to evaluate the proposed method on the CIFAR-100 and STL-10 datasets.
-Experiment results from their work concluded that with curriculum learning the convergence is faster during the beginning of training and improved generalisation is achieved when more difficult tasks are used.
+A CNN with two architectures trained using curriculum learning was used to evaluate the proposed method on the CIFAR-100 [38] and STL-10 datasets [39].
+Experiment results from their work concluded that with curriculum learning the convergence is faster during the beginning of training and improved generalization is achieved when more difficult tasks are used.
 Hacohen et al.
@@
 For the pacing function, three approaches were used with all pacing functions having comparable performance based on experiments done by the authors.
-Experiments done on 6 test cases showed that using a curriculum provided high accuracy and better convergence with the amount of improvement ranging from small to large depending on the test case.
+Experiments done on six test cases showed that using a curriculum provided high accuracy and better convergence with the amount of improvement ranging from small to large depending on the test case.
 Huang et al.
@@
 Zhou et al.
-presented a curriculum learning method known as dynamics-optimized curriculum learning (DoCL) [22].
+[26] presented a curriculum learning method known as dynamics-optimized curriculum learning (DoCL).
 DoCL selected training samples at each step using weighted sampling based on the scores.
-The authors conducted experiments on more than 9 datasets achieving results that significantly improved the performance and efficiency compared to existing curriculum learning methods.
+The authors conducted experiments on more than nine datasets achieving results that significantly improved the performance and efficiency compared to existing curriculum learning methods.
 II-C Data density based curriculum learning
-Our review of literature for curriculum learning has found few papers that utilise the density of the data.
+Our review of literature for curriculum learning has found few papers that utilize the density of the data.
 These works are presented in this section and our work on DDCL adds to this field of research.
-Applying curriculum learning to large data sets has shown promising results.
+Specifically, DDCL addresses the key limitation of other works [25, 34] that require large datasets to utilize data density by effectively operating on small datasets.
+Moreover, clustering in DDCL is used to determine the class centroids rather than the density compared to other works.
 Researchers in [25] presented a method for training deep neural networks on large-scale weakly-supervised web images.
-Their training strategy utilised curriculum learning to effectively handle the large amount of noisy labels and data imbalance during the training process.
+Their training strategy utilized curriculum learning to effectively handle the large amount of noisy labels and data imbalance during the training process.
 They proposed a curriculum created by measuring the complexity of data using the density of its data.
-Experiments using the proposed training strategy resulted in state-of-the-art performance on benchmark data sets that included Webvision, ImageNet, Clothing-1M and Food-101.
+Experiments using the proposed training strategy resulted in state-of-the-art performance on benchmark datasets.
 Choi et al.
@@
 Experiments achieved state-of-the-art performance on the CelebA face attribute dataset and the RAP pedestrian attribute dataset.
-// Variables:
-// Classes: $S$
-// Data size: $N$
-// $N_{s}$ data set for class $s$
-// $|N_{s}|$ data size (cardinality)
-// $h$ bandwidth
-1 foreach $s\in S$ do
-2       ${O_{s}}=\sum_{n=0}^{|N_{s}|}\min_{\mu_{s}\in S}(\parallel x_{n}-\mu_{s}\parallel^{2})$
-3       $s=s\cup O_{s}$
-4       $E_{s}=\varnothing$
-5       foreach $x\in s$ do
-6             $E_{s}=E_{s}\>\cup\parallel{O_{s}-x}\parallel$
-7
-8       end foreach
-9      $D_{s}=\frac{1}{|N_{s}|h}\sum_{i=1}^{|N_{s}|}K(\frac{y-\hat{E}_{i}}{h})$
-10       for $q\in Q^{s}$ do
-11             if $|q|<3$ then
-12                   $q=SMOTE(q)$
-13             end if
-14
-15       end for
-16      if Scoring type normal-$\leftarrow$ “Density” then
-17            go to Algorithm 2
-18       else if Scoring type normal-$\leftarrow$ “Point” then
-19            go to Algorithm 3
-20       end if
-      Result: $T\leftarrow R_{s}$
-21
-22 end foreach
+(a) Density scoring showing the densities of each quantile.
+(b) Point scoring showing samples sorted by the $\hat{E_{s}}$ value.
+Figure 2: Example of DDCL scoring methods on Haberman’s Survival data.
 Algorithm 1 DDCL algorithm
+1:  Variables:
+2:  Classes: $S$
+3:  Data size: $N$
+4:  $N_{s}$ dataset for class $s$
+5:  $|N_{s}|$ data size (cardinality)
+6:  $h$ bandwidth
+7:  for all $s\in S$ do
+8:     ${O_{s}}\leftarrow\sum_{n=0}^{|N_{s}|}\min_{\mu_{s}\in S}(\parallel x_{n}-\mu_{s}\parallel^{2})$
+9:     $s\leftarrow s\cup O_{s}$
+10:     $E_{s}\leftarrow\varnothing$
+11:     for all $x\in s$ do
+12:        $E_{s}\leftarrow E_{s}\cup\parallel{O_{s}-x}\parallel$
+13:     end for
+14:     $D_{s}\leftarrow\frac{1}{|N_{s}|h}\sum_{i=1}^{|N_{s}|}K\left(\frac{y-\hat{E}_{i}}{h}\right)$Kernel Density Estimation}
+15:     for $q\in Q^{s}$ do
+16:        if $|q|<3$ then
+17:           $q\leftarrow SMOTE(q)$
+18:        end if
+19:     end for
+20:     if Scoring type $\leftarrow$ “Density” then
+21:        go to Algorithm 2
+22:     else if Scoring type $\leftarrow$ “Point” then
+23:        go to Algorithm 3
+24:     end if
+25:
+26:     return  $T\leftarrow R_{s}$
+27:  end for
 III Proposed Curriculum Learning Approach
 The curriculum learning approach proposed in this paper is based on the data distribution of a dataset.
-This is referred to as Data Distribution-based Curriculum Learning (DDCL) and involves multiple steps as visualised through Figure 1.
+This is referred to as Data Distribution-based Curriculum Learning (DDCL) and involves multiple steps as visualized through Figure 1.
 It begins with dividing the data into groups based on their target classes.
-Then, the centroids for each group is calculated (Centroid Determination) and used to compute the Euclidean distance between each sample and its centroid.
-The data distribution for every class is determined next and utilised to divide the data for individual classes into quantiles (Q1, Q2, Q3).
-After quantile division, oversampling is optionally performed on quantiles where the number samples may be highly imbalanced compared to other quantiles.
+Then, the centroids for each group are calculated (Centroid Determination) and used to compute the Euclidean distance between each sample and its centroid.
+The data distribution for each class is determined next and utilized to divide the data for individual classes into quantiles (Q1, Q2, Q3).
+After quantile division, oversampling is optionally performed on quantiles where the number of samples may be highly imbalanced compared to other quantiles.
 Lastly, the data is scored with either density or point scoring and rearranged for use in training.
-The DDCL process is further detailed in Algorithm 1.
+The DDCL process is further detailed in Algorithm 1, Algorithm 2 and Algorithm 3.
 In the first step of DDCL, the training data samples are grouped according to their classes ($S$) with the number of groups being equal to the number of unique classes in the data.
@@
 Then, the centroids are used to compute the Euclidean distance from the relevant centroid ($E_{s}$) for each data sample ($x$).
-The third step determines the data distribution for a class ($D_{s}$) by plotting the normalised values of the Euclidean distances.
+The third step determines the data distribution for a class ($D_{s}$) by plotting the normalized values of the Euclidean distances.
 Once the distribution of the training data is known, the data samples are divided into quantiles ($Q^{s}$).
@@
 Finally, the rearranged training data ($T$) is ready to be passed to one of the learning methods for training and evaluation.
-(a) Density scoring
-(b) Point scoring
-Figure 2: DDCL scoring methods
 DDCL consists of two types of scoring methods: a sample density based method and a Euclidean distance based method.
-These are referred to as DDCL (Density) and DDCL (Point) respectively and are detailed in the following sections.
+These are referred to as DDCL-Density and DDCL-Point respectively and are detailed in the following sections.
+TABLE I: Datasets used for evaluating DDCL.
+Dataset Data Instances Attributes Class Type
+Breast Cancer Wisconsin (Diagnostic) 569 31 Binary
+Cancer 457 9 Binary
+Haberman’s Survival 306 3 Binary
+Liver Disorder 345 6 Binary
+Pima Indians Diabetes 768 8 Binary
+New-Thyroid 215 5 Multi-class
+Diabetes 130 [42] 86556 46 Multi-class
 III-A Density based DDCL
-Figure 1(a) illustrates the DDCL (Density) process where the scoring is applied to the quantiles based on the number of samples in the quantile.
-Quantiles with the greatest number of samples (Q3 - 86) are given a higher score while quantiles with the least number of samples (Q1 - 12) are given a lower score.
-Higher scores are considered simple while lower scores are considered difficult.
+Figure 2(a) illustrates the DDCL-Density process where the scoring is applied to the quantiles based on the number of samples in the quantile.
+Quantiles with the greatest number of samples (Q6) are given a higher score whereas quantiles with the least number of samples (Q1) are given a lower score.
+Higher scores are considered simple and lower scores are considered difficult.
 This results in a final quantile order sorted from the highest to lowest density thus defining the curriculum to be used during training.
-[…]
-2 foreach $q\in Q^{s}$ do
-3       $C_{s}=C_{s}\cup|q|$
-4 end foreach
-[…]
-return $R_{s}$
+TABLE II: Neural Network results with and without DDCL.
+Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
+Breast Cancer (Diagnostic) No Curriculum 91.304 100.000 96.232 $\pm$2.442
+DDCL-Density 89.130 100.000 96.594 $\pm$2.673
+DDCL-Point 86.957 100.000 97.319 $\pm$2.953
+Cancer No Curriculum 86.486 100.000 94.955 $\pm$3.993
+DDCL-Density 91.667 100.000 96.852 $\pm$2.658
+DDCL-Point 86.111 100.000 96.574 $\pm$3.100
+Haberman’s Survival No Curriculum 45.833 79.167 65.833 $\pm$8.898
+DDCL-Density 50.000 87.500 69.167 $\pm$10.353
+DDCL-Point 41.667 87.500 67.222 $\pm$9.424
+Liver Disorder No Curriculum 57.143 85.714 69.048 $\pm$6.477
+DDCL-Density 46.429 85.714 68.214 $\pm$9.052
+DDCL-Point 57.143 82.143 69.524 $\pm$6.098
+Pima Indians Diabetes No Curriculum 54.098 86.885 71.803 $\pm$6.658
+DDCL-Density 55.738 85.246 72.732 $\pm$5.933
+DDCL-Point 57.377 80.328 71.803 $\pm$5.460
+New-Thyroid No Curriculum 82.353 100.000 93.137 $\pm$5.280
+DDCL-Density 82.353 100.000 95.490 $\pm$5.409
+DDCL-Point 82.353 100.000 94.118 $\pm$5.683
+Diabetes 130 No Curriculum 53.235 54.694 54.029 $\pm$0.518
+DDCL-Density 56.148 56.798 56.579 $\pm$0.246
+DDCL-Point 54.400 56.278 55.139 $\pm$0.696
+TABLE III: SVM results with and without DDCL.
+Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
+Breast Cancer (Diagnostic) No Curriculum 94.737 99.415 97.700 $\pm$1.188
+DDCL-Density 95.322 99.415 97.758 $\pm$1.025
+DDCL-Point 94.737 99.415 97.719 $\pm$1.239
+Cancer No Curriculum 94.928 99.275 96.715 $\pm$1.134
+DDCL-Density 94.118 99.265 96.985 $\pm$1.145
+DDCL-Point 92.647 99.265 96.225 $\pm$1.562
+Haberman’s Survival No Curriculum 65.217 78.261 72.609 $\pm$4.051
+DDCL-Density 65.217 83.696 73.841 $\pm$4.836
+DDCL-Point 67.391 79.348 73.080 $\pm$3.329
+Liver Disorder No Curriculum 44.231 76.923 63.750 $\pm$7.264
+DDCL-Density 49.038 75.000 66.699 $\pm$5.515
+DDCL-Point 55.769 75.962 68.558 $\pm$4.783
+Pima Indians Diabetes No Curriculum 71.429 80.519 76.046 $\pm$2.500
+DDCL-Density 70.996 82.251 77.128 $\pm$2.991
+DDCL-Point 72.294 82.684 76.508 $\pm$2.514
+New-Thyroid No Curriculum 92.308 100.000 97.128 $\pm$2.472
+DDCL-Density 92.308 100.000 97.282 $\pm$1.892
+DDCL-Point 92.308 100.000 96.872 $\pm$1.563
+Diabetes 130 No Curriculum 64.285 65.487 64.948 $\pm$0.302
+DDCL-Density 64.220 65.457 64.977 $\pm$0.298
+DDCL-Point 64.301 65.469 65.052 $\pm$0.266
+TABLE IV: Random Forest results with and without DDCL.
+Dataset Test Scenario Worst % Best % Average ±$\pm\sigma$
+Breast Cancer (Diagnostic) No Curriculum 94.152 98.246 95.945 $\pm$1.034
+DDCL-Density 92.398 98.246 95.653 $\pm$1.581
+DDCL-Point 91.228 98.246 95.945 $\pm$1.517
+Cancer No Curriculum 92.754 100.000 97.150 $\pm$1.374
+DDCL-Density 94.853 100.000 96.691 $\pm$1.298
+DDCL-Point 94.118 100.000 97.206 $\pm$1.361
+Haberman’s Survival No Curriculum 58.696 76.087 68.261 $\pm$4.100
+DDCL-Density 65.217 79.348 70.145 $\pm$3.825
+DDCL-Point 57.609 77.174 69.746 $\pm$4.367
+Liver Disorder No Curriculum 59.615 78.846 70.417 $\pm$5.104
+DDCL-Density 67.308 77.885 72.917 $\pm$2.709
+DDCL-Point 60.577 78.846 70.064 $\pm$4.077
+Pima Indians Diabetes No Curriculum 71.861 79.654 75.483 $\pm$1.985
+DDCL-Density 70.996 80.087 75.339 $\pm$2.200
+DDCL-Point 65.368 81.818 76.003 $\pm$3.461
+New-Thyroid No Curriculum 95.385 100.000 97.949 $\pm$1.214
+DDCL-Density 92.308 100.000 98.000 $\pm$1.911
+DDCL-Point 95.385 100.000 97.897 $\pm$1.512
+Diabetes 130 No Curriculum 65.830 66.658 66.218 $\pm$0.228
+DDCL-Density 65.626 66.948 66.242 $\pm$0.309
+DDCL-Point 65.542 66.805 66.199 $\pm$0.293
 Algorithm 2 DDCL Density
-Algorithm 2 provides details of the DDCL (Density) scoring process.
+1:  $C_{s}\leftarrow\varnothing$Initialize}
+2:  for all $q\in Q^{s}$ do
+3:     $C_{s}\leftarrow C_{s}\cup|q|$
+4:  end for
+5:  $R_{s}\leftarrow sort(\forall\>C_{s}\mid s\in S)$
+6:
+7:  return  $R_{s}$
+Algorithm 2 provides details of the DDCL-Density scoring process.
 It shows that the process starts with determining the density of each quantile by taking their cardinality, $|q|$.
@@
 III-B Point based DDCL
-On the other hand, Figure 1(b) illustrates how DDCL (Point) assigns a score to the individual data samples in the dataset by utilising their normalised Euclidean distance from centroid rather than examining quantiles.
-In this scoring method, data samples with the shorter Euclidean distances (P2, P5) are given the highest scores while samples with the longer Euclidean distances (P7, Pi) are assigned the lowest scores.
-As with DDCL (Density), higher scores in this point scoring method are considered simple while lower scores are considered difficult.
-Using DDCL (Point) results in a curriculum where the training samples are ordered based on their individual characteristic rather than as a group.
-[…]
-return $R_{s}$
+On the other hand, Figure 2(b) illustrates how DDCL-Point assigns a score to the individual data samples in the dataset by utilizing their normalized Euclidean distance from centroid rather than examining quantiles.
+In this scoring method, data samples with the shorter Euclidean distances (1 and 5) are given the highest scores whereas samples with the longer Euclidean distances (3 and $s$) are assigned the lowest scores.
+As with DDCL-Density, higher scores in this point scoring method are considered simple whereas lower scores are considered difficult.
+Using DDCL-Point results in a curriculum where the training samples are ordered based on their individual characteristic rather than as a group.
 Algorithm 3 DDCL Point
-Algorithm 3 explains the DDCL (Point) scoring method and it shows that the data samples are sorted using the normalised Euclidean distance from centroid, $\hat{E_{s}}$.
+1:  $R_{s}\leftarrow sort(\forall\>\hat{E_{s}}|s\in S)$
+2:
+3:  return  $R_{s}$
+Algorithm 3 explains the DDCL-Point scoring method and it shows that the data samples are sorted using the normalized Euclidean distance from centroid, $\hat{E_{s}}$.
 The sorting is done from the shortest distance to the longest regardless of which quantile a sample is assigned to.
@@
 Next, the fourth lowest $\hat{E_{s}}$ may come from a sample in quantile 3.
-This method of sorting then repeats until $\hat{E_{s}}$ values from all data samples are accounted for with the resulting order of quantiles and samples ($R_{s}$) becoming randomised in contrast to DDCL (Density).
+This method of sorting then repeats until $\hat{E_{s}}$ values from all data samples are accounted for with the resulting order of quantiles and samples ($R_{s}$) becoming randomized in contrast to DDCL-Density.
+Figure 3: Precision-Recall curves for binary classification datasets.
+Figure 4: Confusion matrix for multi-class classification datasets.
 IV Experiments
-DDCL is evaluated by performing classification tasks on 7 datasets obtained from the UCI Machine Learning Repository [35].
-Table I lists the datasets used for the experiments.
+DDCL is evaluated by performing classification tasks on seven datasets obtained from the UCI Machine Learning Repository [43].
+As listed in Table I, the datasets used for the experiments focus on small to medium-sized medical data.
 Three types of classifiers are used for evaluating DDCL; neural network, SVM and random forest.
-[b] Dataset Data Instances Attributes Class Type Breast Cancer Diagnostic 569 31 Binary Cancer 457 9 Binary Haberman 306 3 Binary Liver Disorder 346 6 Binary Pima Indians Diabetes 768 8 Binary New-Thyroid 215 5 Multi-class Diabetes 1301 86556 46 Multi-class
-TABLE I: Datasets used for evaluating DDCL.
-  • 1
-    Preprocessed version found in [36].
+TABLE V: Comparison of DDCL results with other works.
+Dataset Method Year Accuracy %
+Breast Cancer Wisconsin (Diagnostic) SVM [44] 2021 97.20
+SVM (DDCL) 2024 99.42
+Haberman’s Survival Random Forest [45] 2023 74.00
+Random Forest (DDCL) 2024 79.35
+Liver Disorder Neural Network [46] 2023 76.67
+Neural Network (DDCL) 2024 82.14
+Pima Indians Diabetes Random Forest [47] 2022 79.57
+Random Forest [45] 2023 79.00
+Random Forest (DDCL) 2024 81.82
+Diabetes 130 Random Forest [42] 2018 55.97
+Random Forest (DDCL) 2024 66.95
 The neural network architecture used for evaluating DDCL consists of one input layer, a variable number of hidden layers and one output layer.
@@
 Similarly, the number of hidden layers changes based on the dataset and is a hyperparameter that can be tuned.
-Bayesian optimisation is used on the model to tune the hidden layer values.
+Bayesian optimization is used on the model to tune the hidden layer values.
 Finally, the size of the output layer can be configured for binary or multi-class classification.
@@
 The gamma ($\gamma$) value of the RBF kernel scales according to the dataset used.
-A RBF kernel is chosen due to its good generalisation ability and robustness to input noise.
+A RBF kernel is chosen due to its good generalization ability and robustness to input noise.
 The random forest classifier in our evaluation of DDCL uses 100 estimators and a value of 2 features for splitting a node.
@@
 The SVM and random forest classifiers are trained using the training subset of the data and tested on the testing subset.
-For each dataset and classifier, 30 experiments runs are performed using three test scenarios: No Curriculum, DDCL (Density) and DDCL (Point).
+For each dataset, 30 experiments runs are performed for each of the three test scenarios: No Curriculum, DDCL-Density and DDCL-Point.
+The same number of experiment runs is used for all three classifiers.
 The first No Curriculum test scenario takes the input data and uses it to train a neural network, SVM and random forest classifier without altering the order of the training data.
 After training, classification experiments are conducted on the testing subset of data.
-In the second DDCL (Density) test scenario, the input data is processed using the steps outlined in the DDCL method and scored by density.
+In the second DDCL-Density test scenario, the input data is processed using the steps outlined in the DDCL method and scored by density.
 Classification is then performed on the testing subset as stated earlier.
-The DDCL (Point) test scenario processes the input data through DDCL with the point scoring approach being used for the scoring step.
+The DDCL-Point test scenario processes the input data through DDCL with the point scoring approach being used for the scoring step.
 Experiments for classification are likewise performed on the testing subset.
-TABLE II: Neural Network classification results.
-TABLE III: SVM classification results.
-Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
-Breast Cancer Diagnostic No Curriculum 91.304 100.000 96.232 $\pm$2.442
-DDCL (Density) 89.130 100.000 96.594 $\pm$2.673
-DDCL (Point) 86.957 100.000 97.319 $\pm$2.953
-Cancer No Curriculum 86.486 100.000 94.955 $\pm$3.993
-DDCL (Density) 91.667 100.000 96.852 $\pm$2.658
-DDCL (Point) 86.111 100.000 96.574 $\pm$3.100
-Haberman No Curriculum 45.833 79.167 65.833 $\pm$8.898
-DDCL (Density) 50.000 87.500 69.167 $\pm$10.353
-DDCL (Point) 41.667 87.500 67.222 $\pm$9.424
-Liver Disorder No Curriculum 57.143 85.714 69.048 $\pm$6.477
-DDCL (Density) 46.429 85.714 68.214 $\pm$9.052
-DDCL (Point) 57.143 82.143 69.524 $\pm$6.098
-Pima Indians Diabetes No Curriculum 54.098 86.885 71.803 $\pm$6.658
-DDCL (Density) 55.738 85.246 72.732 $\pm$5.933
-DDCL (Point) 57.377 80.328 71.803 $\pm$5.460
-New-Thyroid No Curriculum 82.353 100.000 93.137 $\pm$5.280
-DDCL (Density) 82.353 100.000 95.490 $\pm$5.409
-DDCL (Point) 82.353 100.000 94.118 $\pm$5.683
-Diabetes 130 No Curriculum 53.235 54.694 54.029 $\pm$0.518
-DDCL (Density) 56.148 56.798 56.579 $\pm$0.246
-DDCL (Point) 54.400 56.278 55.139 $\pm$0.696
-Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
-Breast Cancer Diagnostic No Curriculum 94.737 99.415 97.700 $\pm$1.188
-DDCL (Density) 95.322 99.415 97.758 $\pm$1.025
-DDCL (Point) 94.737 99.415 97.719 $\pm$1.239
-Cancer No Curriculum 94.928 99.275 96.715 $\pm$1.134
-DDCL (Density) 94.118 99.265 96.985 $\pm$1.145
-DDCL (Point) 92.647 99.265 96.225 $\pm$1.562
-Haberman No Curriculum 65.217 78.261 72.609 $\pm$4.051
-DDCL (Density) 65.217 83.696 73.841 $\pm$4.836
-DDCL (Point) 67.391 79.348 73.080 $\pm$3.329
-Liver Disorder No Curriculum 44.231 76.923 63.750 $\pm$7.264
-DDCL (Density) 49.038 75.000 66.699 $\pm$5.515
-DDCL (Point) 55.769 75.962 68.558 $\pm$4.783
-Pima Indians Diabetes No Curriculum 71.429 80.519 76.046 $\pm$2.500
-DDCL (Density) 70.996 82.251 77.128 $\pm$2.991
-DDCL (Point) 72.294 82.684 76.508 $\pm$2.514
-New-Thyroid No Curriculum 92.308 100.000 97.128 $\pm$2.472
-DDCL (Density) 92.308 100.000 97.282 $\pm$1.892
-DDCL (Point) 92.308 100.000 96.872 $\pm$1.563
-Diabetes 130 No Curriculum 64.285 65.487 64.948 $\pm$0.302
-DDCL (Density) 64.220 65.457 64.977 $\pm$0.298
-DDCL (Point) 64.301 65.469 65.052 $\pm$0.266
-TABLE III: SVM classification results.
-TABLE IV: Random Forest classification results.
-Dataset Test Scenario Worst % Best % Average $\pm$±Std. deviation
-Breast Cancer Diagnostic No Curriculum 94.152 98.246 95.945 $\pm$1.034
-DDCL (Density) 92.398 98.246 95.653 $\pm$1.581
-DDCL (Point) 91.228 98.246 95.945 $\pm$1.517
-Cancer No Curriculum 92.754 100.000 97.150 $\pm$1.374
-DDCL (Density) 94.853 100.000 96.691 $\pm$1.298
-DDCL (Point) 94.118 100.000 97.206 $\pm$1.361
-Haberman No Curriculum 58.696 76.087 68.261 $\pm$4.100
-DDCL (Density) 65.217 79.348 70.145 $\pm$3.825
-DDCL (Point) 57.609 77.174 69.746 $\pm$4.367
-Liver Disorder No Curriculum 59.615 78.846 70.417 $\pm$5.104
-DDCL (Density) 67.308 77.885 72.917 $\pm$2.709
-DDCL (Point) 60.577 78.846 70.064 $\pm$4.077
-Pima Indians Diabetes No Curriculum 71.861 79.654 75.483 $\pm$1.985
-DDCL (Density) 70.996 80.087 75.339 $\pm$2.200
-DDCL (Point) 65.368 81.818 76.003 $\pm$3.461
-New-Thyroid No Curriculum 95.385 100.000 97.949 $\pm$1.214
-DDCL (Density) 92.308 100.000 98.000 $\pm$1.911
-DDCL (Point) 95.385 100.000 97.897 $\pm$1.512
-Diabetes 130 No Curriculum 65.830 66.658 66.218 $\pm$0.228
-DDCL (Density) 65.626 66.948 66.242 $\pm$0.309
-DDCL (Point) 65.542 66.805 66.199 $\pm$0.293
-The results of all three test scenarios using neural networks, SVM and random forest classifiers are given in Table III, Table III and Table IV respectively.
-Figure 3: Precision-Recall curves for binary classification datasets and confusion matrix for multi-class datasets.
+Figure 5: Error loss per epoch for neural network classifier.
+The results for neural networks, SVM and random forest classifiers are given in Table II, Table III and Table IV respectively.
 V Discussion
 Observations on the learning methods show that the classification performance across the datasets is always increased whenever DDCL is applied.
-This is true regardless of the learning method used as shown in Table III, Table III and Table IV.
-Figure 4: Error loss per epoch for neural network classifier.
-Using the neural network classifier provides the highest best results for the Haberman, Liver Disorder and Pima Indians Diabetes datasets while the SVM method achieves the highest average results on the Breast Cancer Diagnostic, Haberman and Pima Indians Diabetes data.
+This is true regardless of the learning method used as shown in Table II, Table III and Table IV.
+Figure 6: Error loss per epoch with batch gradient descent for neural network classifier.
+Using the neural network classifier provides the highest best results for the Haberman’s Survival, Liver Disorder and Pima Indians Diabetes datasets whereas the SVM method achieves the highest average results on the Breast Cancer (Diagnostic), Haberman’s Survival and Pima Indians Diabetes data.
 The random forest based classifier yields the highest average result for the Cancer, Liver Disorder, New-Thyroid and Diabetes 130 datasets.
-Notably, using DDCL with SVM and random forest approaches results in significant increases in average accuracy for Haberman and Liver Disorder data respectively compared to other learning methods.
+Notably, using DDCL with SVM and random forest approaches results in significant increases in average accuracy for Haberman’s Survival and Liver Disorder data respectively compared to other learning methods.
 Figure 3 shows the precision-recall curves for the best result of each binary classification dataset when using the neural network method.
-For multi-class data, the confusion matrix is provided in place of the precision-recall curves.
-The highest average result for Breast Cancer Diagnostic was achieved using SVM with density-based DDCL while the largest improvement (1.087%) was using the neural network with point-based DDCL.
+For multi-class data, the confusion matrix is provided in place of the precision-recall curves by Figure 4.
+Table V shows a comparison of results obtained using DDCL compared to approaches used by other state-of-the-art works.
+It demonstrates DDCL’s ability to achieve superior results compared to the same type of classifier trained using a different algorithm.
+In particular it outperforms the balanced stratified method used in [45] on the challenging Haberman’s Survival and Pima Indians Diabetes datasets.
+Furthermore, DDCL provides greater generalization capability and performance with neural networks on the Liver Disorder data compared to the standard approach used in [46].
+The highest average result for Breast Cancer (Diagnostic) was achieved using SVM with density-based DDCL while the largest improvement (1.087%) was using the neural network with point-based DDCL.
 For the Cancer dataset, using random forest with point-based DDCL provided the highest average accuracy and using the neural network with density-based DDCL gave the most improved accuracy (1.897%).
-Using SVM with density-based DDCL resulted in the highest average accuracy when evaluated on the Haberman data.
+Using SVM with density-based DDCL resulted in the highest average accuracy when evaluated on the Haberman’s Survival data.
 The largest improvement for this dataset (3.334%) was observed when using density-based DDCL with the neural network.
-The highest average result for Liver Disorder was achieved using random forest with density-based DDCL while the SVM gave the largest average accuracy improvement at 4.808%.
+The highest average result for Liver Disorder was achieved using random forest with density-based DDCL whereas the SVM gave the largest average accuracy improvement at 4.808%.
 For the Pima Indians Diabetes dataset, the highest average accuracy was obtained using SVM with density-based DDCL which was also the most improved accuracy (1.082%).
-In multi-class classification, using random forest with density-based DDCL resulted in the highest average accuracy on the New-Thyroid data and the largest improvement for this dataset came from density-based DDCL applied to the neural network.
-For the Diabetes 130 data, density-based DDCL with random forest provided the highest average accuracy while the largest improvement in accuracy (2.550%) was obtained through the neural network classifier utilising density-based DDCL.
+In multi-class classification, using random forest with density-based DDCL resulted in the highest average accuracy on the New-Thyroid data and the largest improvement for this dataset (2.353%) came from density-based DDCL applied to the neural network.
+For the Diabetes 130 data, density-based DDCL with random forest provided the highest average accuracy while the largest improvement in accuracy (2.550%) was obtained through the neural network classifier utilizing density-based DDCL.
 The experiment results demonstrate that using DDCL to order the training data for a given learning method leads to improvements in classification accuracy.
@@
 Figure 5 shows the error loss per epoch plots when using the neural network classifier for each dataset except for Diabetes 130 which is excluded since its average accuracy is not significant [42].
-The plots show that using DDCL leads to better reduction in the error loss.
+The plot for Breast Cancer (Diagnostic) shows a minimal change in the DDCL error loss trends compared to the no curriculum method whereas the Cancer dataset plot presents a reduction in the error loss when using DDCL.
+For the remaining datasets, the use of DDCL shows a significant and quick reduction in the error loss.
 In addition, using DDCL to order the training data results in faster convergence.
@@
 Figure 6 shows the error loss plots for each dataset.
-It can be seen that the overall losses for DDCL (Density) and DDCL (Point) are reduced faster compared to the no curriculum approach thus leading to faster convergence towards the minimum.
-This is consistent with the findings of the error loss against epoch plots in Figure 4 where the error loss towards the end is lower for DDCL (Density) and DDCL (Point).
-Figure 5: Error loss per epoch with batch gradient descent for neural network classifier.
+It can be seen that the overall losses for DDCL-Density and DDCL-Point are reduced faster compared to the no curriculum approach thus leading to faster convergence towards the minimum.
+This is consistent with the findings of the error loss against epoch plots in Figure 5 where the error loss towards the end is lower for DDCL-Density and DDCL-Point.
 VI Conclusion
@@
 It used the data distribution of a dataset to build a curriculum based on the order of samples.
-Two types of scoring methods known as DDCL (Density) and DDCL (Point) were used to score samples thus determining their training order.
-DDCL (Density) used the sample density to assign scores while DDCL (Point) utilised the Euclidean distance for scoring.
-Experiments were conducted on multiple datasets using a neural network, SVM and random forest classifier to evaluate DDCL.
-Results showed that while performance varied across the datasets depending on the classifier used, the application of DDCL resulted in significant increases in the average accuracy for most datasets compared to not using any curriculum.
+Two types of scoring methods known as DDCL-Density and DDCL-Point were used to score samples thus determining their training order.
+DDCL-Density used the sample density to assign scores whereas DDCL-Point utilized the Euclidean distance for scoring.
+Experiments were conducted on multiple medical datasets using a neural network, SVM and random forest classifier to evaluate DDCL.
+Results showed that even though performance varied across the datasets between the classifiers used, the application of DDCL resulted in accuracy increases ranging from 2% to 10% for all datasets compared to other state-of-the-art methods.
 Furthermore, analysis of the error losses for five training epochs using batch gradient descent reveals that convergence is faster when using DDCL with either of the scoring methods over the no curriculum method.
@@
 Future work will explore the creation of additional scoring methods to determine their impact on the training performance and investigate the application of ensemble learning using each scoring method.
+By introducing these new scoring methods, we aim to assess their effectiveness in comparison to other curriculum approaches such as SPL and Reverse Curriculum Learning (RCL).
 In addition, self-paced learning concepts will be incorporated into DDCL in order to dynamically determine the curriculum based on feedback from the learner.
-The proposed approach can also be adapted to image and text processing applications.
+This integration will allow the model to adjust its learning pace according to its current performance, potentially leading to a more robust and adaptive training process.
+Furthermore, the algorithms of the proposed DDCL’s scoring methods suggests versatility which can be adapted beyond its initial application to structured datasets.
+Specifically, the proposed approach can also be adapted to image and text processing applications where the inherent complexity and variability of data types present unique challenges.
+Nomenclature
+Symbol Description
+$S$ Grouped data samples for all classes
+$s$ Data for a specific class
+${O_{s}}$ Centroids for a specific class
+$E_{s}$ Euclidean distances from centroid (dimensionless)
+$D_{s}$ Data distribution for a class
+$Q^{s}$ Quantiles for the whole dataset
+$q$ A single quantile in a dataset
+$R_{s}$ Scored training data
+$T$ Curriculum prepared training data
+$C_{s}$ Densities (count) for all quantiles
+$\hat{E_{s}}$ Normalized Euclidean distances from centroid